OpenStudy (anonymous):

In a study of 205 adults, the mean heart rate was 75 beats per minute. Assume the population of heart rates is known to be approximately normal with a standard deviation of 8 beats per minute. What is the 95% confidence interval for the mean beats per minute?

OpenStudy (anonymous):

73.9 – 76.1 73.7 – 76.1 73.9 – 76.3 70.9 – 73.3

OpenStudy (anonymous):

@ClarissaMorgenstern canu you help me

OpenStudy (anonymous):

@RadEn

OpenStudy (kropot72):

Do you know the formula for this calculation? If not, I can post it and help you.

OpenStudy (anonymous):

no

OpenStudy (anonymous):

@kropot72 can u help ?

OpenStudy (kropot72):

The confidence interval for the mean is given by: \[\bar {x}-1.96 \frac{\sigma}{\sqrt{n}}<\mu<\bar {x}+1.96 \frac{\sigma}{\sqrt{n}}\] Substituting the given values we get: \[75-1.96 \frac{8}{\sqrt{205}}<\mu<75+1.96 \frac{8}{\sqrt{205}}\] Can you now do the calculation?

OpenStudy (anonymous):

i can try but i dont even know what im looking at

OpenStudy (kropot72):

First, do this calculation: \[1.96\times \frac{8}{\sqrt{205}}=you\ can\ calculate?\]

OpenStudy (anonymous):

im trying

OpenStudy (anonymous):

1.095138703767

OpenStudy (anonymous):

@kropot72

OpenStudy (anonymous):

@Valpey

OpenStudy (anonymous):

i got a

OpenStudy (kropot72):

Good, you have found the standard error of the mean. This can be rounded up to equal 1.1. So the 95% confidence interval for the mean beats per minute is given by: (75 - 1.1) to (75 + 1.1). Can you now do the subtraction and addition to find the answer?

OpenStudy (anonymous):

A spa owner gives a survey to all of her customers asking them to rate the quality of the service they received. She then keeps track of how many customers return to the spa for additional services during the next six months. Last year, the results showed that of the customers who reported high quality service, 15% returned for additional services. What conclusion can be drawn from this study?

OpenStudy (anonymous):

i got A as the answer

OpenStudy (kropot72):

Yes, A is the correct answer :)

OpenStudy (anonymous):

can u help me with the other one

OpenStudy (kropot72):

Sorry, I have to log out now. Please post the other one as a new question.

OpenStudy (anonymous):

Thankyou guys smm.... for the first the 73.9 – 76.1 was correct